# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v2

@@ 5.1 Proof of Theorem 1
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\tilde{\eta}^{k}_{H}\equiv 0$ for all $k\in[K]$. Summing (8) over $\{h,h+1,\cdots,H-1\}$, the terms telescope, yielding
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (9)
Next, the claim the following inequality holds:
$\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}$
$\qquad\leq\,\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (10)
If $a_{h}^{k}\in\mathcal{A}_{h}^{*}(s_{h}^{k})$, the inequality holds trivially. Suppose that $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Note that,
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $=(V^{k}_{h}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq(V^{*}_{h}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$ (by $\mathbbm{1}\{\mathcal{G}^{k}\}$ and $V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}$ ($a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$)
which implies $\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}=\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ given $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Now, we have the following cases:
  1. 1.
    If $\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})\geq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$, then
@@
    $\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ $=\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
    $\leq-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$ (by (9))
  2. 2.
    If $2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))>-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$, then
    $\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}\leq\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $\leq 4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})),$ (by (9))
    which implies
    $\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)=4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})).$
Substituting (10) into (7) yields
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ $\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$
$\qquad\qquad-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathrm{clip}\left(4Hb^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$ (by (4))
$\qquad\qquad-2\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
$\qquad\qquad-\sum_{h=0}^{H-1}\left(2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k})\right)\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
@@
$=\,\left[\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (11)
where the sequence $\{\bar{\eta}_{h}^{k}\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\bar{\eta}^{k}_{h+1}=2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k}),$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$, and satisfies $|\bar{\eta}^{k}_{h+1}|\leq 5H^{2}$.
Next, we bound the first term in (11). We have
@@
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$ $\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
@@
$\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n_{h}^{k}(s_{h}^{k},a_{h}^{k})\}}}\right)$
$\leq H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\infty}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$=H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\bar{n}}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$\leq 8H^{4}SA\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sqrt{\bar{n}}$
$\leq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}}\eqqcolon m_{K}^{\mathrm{mb}}$ (12)
where
$\bar{n}=\left\lfloor\frac{16H^{4}(0.5S\ln 2+\mu K^{\alpha})}{{\mathrm{gap}^{*}}^{2}}\right\rfloor.$
@@
$\qquad\qquad\leq\mathbb{P}\left(\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\geq x-\lceil K^{\gamma}\rceil H\right)$
$\qquad\qquad\leq\mathbb{P}\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\geq x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}}\right)$ (since $x>\lceil K^{\gamma}\rceil H$)
$\qquad\qquad\leq\exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}})^{2}}{50H^{5}K}\right).$
where the last inequality follows by applying Azuma–Hoeffding.
Proof of Proposition 1.

### v3

@@ 6.1 Proof of Theorem 1
is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\tilde{M}^{k}_{H}\equiv 0$ for all $k\in[K]$. Summing (14) over $\{h,h+1,\cdots,H-1\}$, the terms telescope, yielding
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{M}_{l+1}^{k}(s_{l+1}^{k}).$ (15)
Notice that the action taken at horizon $\upsilon^{k}$ is suboptimal. Hence, we have
$\tilde{\Delta}_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $=\left(V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq\left(V_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}$ (by (12))
$\geq\left(V_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k})-Q_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k},a_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}$ (by $V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})=Q_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k},a_{\upsilon^{k}}^{k})\leq Q_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k},a_{\upsilon^{k}}^{k})$ )
$=\mathrm{gap}_{\upsilon^{k}}(s_{\upsilon^{k}}^{k},a_{\upsilon^{k}}^{k})$
$\geq\mathrm{gap}^{*},$
@@
which implies that $\tilde{\Delta}_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}=\mathrm{clip}\left(\tilde{\Delta}_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}$. Substituting this into (13) yields
@@
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ $\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\mathrm{clip}\left(\tilde{\Delta}_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
@@
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\mathrm{clip}\left(2\sum_{l=\upsilon^{k}}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))-\sum_{l=\upsilon^{k}}^{H-1}\tilde{M}_{l+1}^{k}(s_{l+1}^{k})\right)-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (15))
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\mathrm{clip}\left(4\sum_{l=\upsilon^{k}}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)-2\sum_{l=\upsilon^{k}}^{H-1}\tilde{M}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by Lemma 4)
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\mathrm{clip}\left(4\sum_{l=\upsilon^{k}}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)-\sum_{l=0}^{H-1}\bar{M}_{l+1}^{k}(s_{l+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (16)
where the sequence $\{\bar{M}^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\bar{M}_{h}^{k}(s_{h}^{k})\coloneqq M_{h}^{k}(s_{h}^{k})\mathbbm{1}\{h\leq\upsilon^{k}\}+2\tilde{M}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{h>\upsilon^{k}\}.$
is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. Crucially, because $\upsilon^{k}$ is the step of the first suboptimal action, the event $\{h\leq\upsilon^{k}\}$ depends only on whether actions $a_{0}^{k},\dots,a_{h-1}^{k}$ are optimal, meaning both $\mathbbm{1}\{h\leq\upsilon^{k}\}$ and $\mathbbm{1}\{h>\upsilon^{k}\}$ are $\mathcal{F}_{h-1}^{k}$-measurable. Therefore, $\bar{M}_{h}^{k}$ is a valid martingale difference sequence with respect to $\mathcal{F}_{h-1}^{k}$. Also, we have $\mathbbm{1}\{\mathcal{G}^{k}\}|\bar{M}_{h}^{k}(s_{h}^{k})|\leq 2(H-h)$. Next, we bound the first term in (16). We have
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathrm{clip}\left(4\sum_{l=\upsilon^{k}}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$ $\leq\sum_{k=0}^{K-1}\mathrm{clip}\left(4\sum_{l=0}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$
$=4\sum_{k=0}^{K-1}\sum_{n=1}^{N}w^{k}(n)\sum_{l=0}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$
$=4\sum_{n=1}^{N}\sum_{k=0}^{K-1}w^{k}(n)\sum_{l=0}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$ (17)
where $w^{k}(n)\coloneqq\mathbbm{1}\{4\sum_{l=0}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\in[2^{n-1}\mathrm{gap}^{*},2^{n}\mathrm{gap}^{*})\}$, $N=\lceil\log_{2}\left(4Hb_{\max}/\mathrm{gap}^{*}\right)\rceil$, and $b_{\max}=H\sqrt{0.5S\ln 2+\mu K^{\alpha}}$. For a fixed $n\in[N+1]$, we have
$\sum_{k=0}^{K-1}w^{k}(n)\sum_{l=0}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$ $=\sum_{l=0}^{H-1}\sum_{k=0}^{K-1}w^{k}(n)b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$
$=\sum_{l=0}^{H-1}\sum_{(s,a)}\,\sum_{\begin{subarray}{c}k=0\\ (s_{l}^{k},a_{l}^{k})=(s,a)\end{subarray}}^{K}w^{k}(n)b^{k}_{l}(n_{l}^{k}(s,a))$
$\leq\sum_{l=0}^{H-1}\sum_{(s,a)}\,\sum_{i=1}^{n_{l}^{K}(s,a)}w^{\tau(s,a,i)}(n)\frac{b_{\max}}{\sqrt{i}}$
$\leq\sum_{l=0}^{H-1}\sum_{(s,a)}\,\sum_{i=1}^{C_{l}{(s,a;n)}}\frac{b_{\max}}{\sqrt{i}}$
$\leq 2b_{\max}\sum_{l=0}^{H-1}\sum_{(s,a)}\sqrt{C_{l}{(s,a;n)}}$
$\leq 2b_{\max}\sqrt{HSAC(n)}$ (18)
where the last inequality follows by Cauchy-Schwartz and $C(n)=\sum_{(s,a)}\sum_{l=0}^{H-1}C_{l}{(s,a;n)}$, $C_{l}{(s,a;n)}=\sum_{i=1}^{n_{l}^{K}(s,a)}w^{\tau(s,a,i)}(n)$, and $\tau(s,a,i)$ denotes the episode index in which action $(s,a)$ at horizon $l$ is taken for the $i$-th time. On the other hand, we have the following lower bound:
$\sum_{k=0}^{K-1}w^{k}(n)\sum_{l=0}^{H-1}4b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\geq 2^{n-1}\mathrm{gap}^{*}C(n).$ (19)
Hence, combining (18) and (19), we have
$C(n)\leq\frac{b^{2}_{max}HSA}{4^{n-4}(\mathrm{gap}^{*})^{2}}$
Hence, (17) is bounded as follows:
@@
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathrm{clip}\left(4\sum_{l=\upsilon^{k}}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$ $\leq\sum_{n=1}^{N}\frac{8b^{2}_{\max}HSA}{2^{n-4}\mathrm{gap}^{*}}\leq\frac{128H^{3}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}}\eqqcolon m_{K}^{\mathrm{mb}}$ (20)
@@
$\qquad\qquad\leq\mathbb{P}\left(\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{l=0}^{H-1}\bar{M}_{h+1}^{k}\right)\geq x-\lceil K^{\gamma}\rceil H\right)$
$\qquad\qquad\leq\mathbb{P}\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{l=0}^{H-1}\bar{M}_{h+1}^{k}\mathbbm{1}\{\mathcal{G}^{k}\}\geq x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}}\right)$ (since $x>\lceil K^{\gamma}\rceil H$)
$\qquad\qquad\leq\exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}})^{2}}{4H(H+1)(2H+1)K/3}\right).$
where the last inequality follows by applying Azuma–Hoeffding. $\square$
6.2 Proof of Proposition 1